# MVMS-RCN: A Dual-Domain Unfolding CT Reconstruction with Multi-sparse-view and Multi-scale Refinement-correction

### v1

@@ IV-B4 Ablation study
TABLE V: Quantitative sparse-view CT reconstruction performance comparisons of different methods. The best and second best results are highlighted in Bold font and underlined ones, respectively.
@@
IV-B5 Module sharing configurations
To demonstrate the superiority of the proposed multi-scale geometric correction module $\mathcal{D}$ with shared network parameter configurations among different stages, we conduct two variants of MVMS-RCN with different shared settings between stages. Table IV lists the mean PSNR and SSIM comparisons with different shared settings of the proposed MVMS-RCN. The default shared variant (a) performs better than the unshared variant (b), indicating that the proposed multi-view projection refinement module and the multi-scale geometric correction module are powerful enough to reduce the learning cost of the multi-scale geometric correction module $\mathcal{D}$ by sharing the learning parameters. We adopt the default shared variant (a) for the following experiments.
We attribute the superiority of our method to three factors. Firstly, our approach offers a dual-domain unfolding refinement-correction framework that can refine full-sparse-view projection errors and distill multi-scale geometric corrections to reconstruct sparse-view CT. Secondly, the projection error refinement from sparse-view and full-view can effectively extract the potential projection data. Lastly, a multi-scale geometric correction module inspired by the multigrid scheme can effectively correct the geometric error under multi-scale and multi-resolution.
@@ IV-C Comparison to the State-of-the-art Methods
To further demonstrate the superiority of our approaches, we conduct a group of studies to evaluate the reconstruction performance of ten compared methods on sparse-view CT. We compare our MVMS-RCN with traditional approaches such as FBP and FISTA-TV [60]), as well as DL techniques such as RED-CNN [19], FBPConvNet [20], DU-GAN [26] and Uformer [61]. Additionally, we compare to deep unfolding methods including PD-Net [33], ISTA-Net [34], ISTA-Net+ [34], FISTA-Net [37], AMP-Net-K [39] and Nest-DGIL [41]. PD-Net stands out as a tried-and-true approach for dual-domain reconstruction. Likewise, other deep unfolding methods also perform corrections in both the image domain and the projection domain. The FBP with the ”Ram-Lak” filter, which was implemented with the Iradon transform [58], is utilized as the initial step for the proposed methods and other compared methods. The maximum number of iterations for FISTA-TV is set to 100, and the regularization parameter is tuned to its optimal value. As reported in [37], the number of stages for PD-Net, ISTA-Net, ISTA-Net+, FISTA-Net, AMP-Net-K, Nest-DGIL and our approach are all configured as $7$. All the compared methods are trained with the same training dataset and the same methods with the corresponding works.
Figure 3: The axial reconstruction results from different methods for fan-beam projection with 32 views. (a1),(a3) reference image, (b1)-(b4) FBP, (c1)-(c4) RED-CNN, (d1)-(d4) FBPConvNet, (e1)-(e4) Uformer, (f1)-(f4) PD-Net, (g1)-(g4) ISTA-Net, (h1)-(h4) ISTA-Net+, (i1)-(i4) FISTA-Net, (j1)-(j4) AMP-Net-K, (k1)-(k4) Nest-DGIL, (l1)-(l4) MVMS-RCN. The display windows are [-1150, 350] HU and [-160, 240] HU, respectively.
Figure 4: The axial reconstruction results from different methods for parallel-beam projection with 60 views. (a1)-(a2) reference image, (b1)-(b2) FBP, (c1)-(c2) RED-CNN, (d1)-(d2) FBPConvNet, (e1)-(e2) Uformer, (f1)-(f2) PD-Net, (g1)-(g2) ISTA-Net, (h1)-(h2) ISTA-Net+, (i1)-(i2) FISTA-Net, (j1)-(j2) AMP-Net-K, (k1)-(k2) Nest-DGIL, (l1)-(l2) MVMS-RCN. The display windows are [-1150, 350] HU and [-160, 240] HU, respectively.
@@ IV-C1 Quantitative evaluation
Table V shows the mean PSNR and SSIM values of the reconstruction results of the compared methods with different projection views of the downsampling. It is evident that the two bottom rows have the best results. All of the DL methods and deep unfolding methods have somewhat competitive reconstruction results for fan-beam and parallel-beam projection data; however, our methods clearly outperform the other compared methods in all reconstruction tasks from different down-sampling projections. In particular for fan-beam projection, the proposed MVMS-RCN have a 4.49 dB improvement with 16 views, a 3.20 dB improvement with 32 views, and an average 1.96 dB improvement on five sparse view CT reconstruction tasks. Additionally, our methods have an average 0.39 dB improvement for parallel-beam projection.
We present the boxplots of the Root Mean Square Error of Hounsfield Unit (RMSE (HU)) values for the fan-beam projection with 32 views and the parallel beam projection with 60 views in Fig. 2 to further demonstrate the superiority of our methods. Our methods are able to strike a balance between the use of prior knowledge from models and the powerful learning ability of DL, resulting in excellent generalizability. Furthermore, our methods outperform the compared methods in terms of reconstruction performance.
IV-C2 Qualitative evaluation
Fig. 3 shows the results of CT reconstruction using different methods in the fan-beam projection with 32 views. We can see that the FBP reconstruction has a lot of noise and artifacts. ISTA-Net, ISTA-Net+, FISTA-Net and Nest-DGIL can reduce some of the noise and streaking artifacts, but details are lost. RED-CNN, FBPConvNet and Uformer can remove streaking artifacts effectively, but some small structures may be smoothed out. PD-Net can remove streaking artifacts well and reconstruct high-quality CT images, but the details and texture information are not completely preserved. Our methods (l1)-(l4) have the best performance in terms of removal of noise artifacts and preservation of detail (e.g., tiny blood vessels and bronchi).
TABLE VI: Quantitative reconstruction comparisons (PSNR) of untrained sparse-view projections. The best and second best results are highlighted in Bold font and underlined ones, respectively.
@@ IV-D Blind Reconstruction of Untrained Sparse-view Projections
Table VI presents the blind reconstruction quantitative comparison of untrained sparse-view projections. All compared methods train five models for fan-beam projection (16, 32, 64, 128 and 256 views) and parallel-beam projection (30, 60, 90, 120 and 180 views), while the proposed framework trains a single model for the five sparse-view CT reconstruction tasks. The compared methods cannot directly reconstruct CT images from the untrained sparse-view projections, so their PSNR results are tested by the two closest trained sparse-view models. Our framework outperforms the other methods for fan-beam projection, with an average improvement of 2.51 dB. Specifically, it improves at least 5.25 dB, 2.97 dB, 0.87 dB, and 0.97 dB for 24 views, 48 views, 96 views, and 192 views, respectively. For parallel-beam projection, our framework improves 1.28 dB, 0.21 dB, 0.16 dB, and 0.27 dB for 45 views, 75 views, 105 views, and 150 views, respectively, with an average 0.48 dB improvement. It is noteworthy that our method can use the single multi-sparse-view model to produce more satisfactory results with any untrained sparse-view projections than the other compared methods.
@@ IV-E Plug-and-Play Sparse-view CT Reconstruction
Despite being learned end-to-end, the proposed MVMS-RCN can serve as a plug-and-play denoiser [62] for sparse-view CT reconstruction due to the shared multi-scale geometric correction module $\mathcal{D}$. The convergence curves of PSNR for validation and test sets in relation to the number of stages are illustrated in Fig. 5. Although our MVMS-RCN sets the stage number to 7 as [37], we can achieve near-optimal reconstruction results when the stages are in the range of 6 to 10 during plug-and-play iterations. It fully demonstrates that our method possesses the potential to be applied to plug-and-play reconstruction.
Figure 5: Convergence curves of PSNR for validation and test sets with respect to number of stages.
@@ V Conclusion
In this study, we propose a novel interpretable dual-domain framework that can flexibly deal with multi-sparse-view CT imaging in a single model. The proposed framework integrates the theoretical advantages of model-based methods and the superior reconstruction performance of DL-based methods, providing the generalizability of DL. It can give us a new way to explore the potential projection information, a new perspective on designing an interpretable network, and a flexible model for multi-sparse-view CT reconstruction. Experiments show that the proposed framework outperforms other state-of-the-art methods in terms of qualitative and quantitative evaluaions.

### v2

@@ IV-B5 Ablation study
TABLE VI: Quantitative sparse-view CT reconstruction performance comparisons of different methods. The best and second best results are highlighted in Bold font and underlined ones, respectively.
@@
IV-B6 Module sharing configurations
To demonstrate the superiority of the proposed multi-scale geometric correction module $\mathcal{D}$ with shared network parameter configurations among different stages, we conduct two variants of MVMS-RCN with different shared settings between stages. Table V lists the average PSNR comparisons with different shared settings of the proposed MVMS-RCN. The default shared variant (a) performs better than the unshared variant (b), indicating that the proposed multi-view projection refinement module and the multi-scale geometric correction module are powerful enough to reduce the learning cost of the multi-scale geometric correction module $\mathcal{D}$ by sharing the learning parameters. We adopt the default shared variant (a) for the following experiments.
We attribute the superiority of our method to three factors. Firstly, our approach offers a dual-domain refinement-correction framework that can refine full-sparse-view projection errors and distill multi-scale geometric corrections to reconstruct sparse-view CT. Secondly, the projection error refinement from sparse-view and full-view can effectively extract the potential projection information. Lastly, a multi-scale geometric correction module inspired by the multigrid scheme can effectively correct the geometric error under multi-scale and multi-resolution.
Figure 3: The axial reconstruction results (’1’ and ’3’) and corresponding residuals (’2’ and ’4’) from different methods for fan-beam projection with 32 views. (a1),(a3) reference image, (b1)-(b4) FBP, (c1)-(c4) RED-CNN, (d1)-(d4) FBPConvNet, (e1)-(e4) Uformer, (f1)-(f4) PD-Net, (g1)-(g4) ISTA-Net, (h1)-(h4) ISTA-Net+, (i1)-(i4) FISTA-Net, (j1)-(j4) Nest-DGIL, (k1)-(k4) DPIR, (l1)-(l4) MVMS-RCN. The display windows are [-1150, 350] HU and [-160, 240] HU, respectively.
Figure 4: The axial reconstruction results from different methods for parallel-beam projection with 60 views. (a1)-(a2) reference image, (b1)-(b2) FBP, (c1)-(c2) RED-CNN, (d1)-(d2) FBPConvNet, (e1)-(e2) Uformer, (f1)-(f2) PD-Net, (g1)-(g2) ISTA-Net, (h1)-(h2) ISTA-Net+, (i1)-(i2) FISTA-Net, (j1)-(j2) Nest-DGIL, (k1)-(k2) DPIR, (l1)-(l2) MVMS-RCN. The display windows are [-1150, 350] HU and [-160, 240] HU, respectively.
@@ IV-C Comparison to the State-of-the-art Methods
To further demonstrate the superiority of our approach, we conduct a group of studies to evaluate the reconstruction performance of fourteen compared methods on sparse-view CT. We compare our MVMS-RCN with traditional approaches such as FBP and FISTA-TV [64]), as well as DL techniques such as RED-CNN [19], FBPConvNet [20], DU-GAN [26], Uformer [65] and DRUNet [45]. Furthermore, we compare with deep unfolding methods, including PD-Net [33], ISTA-Net [34], ISTA-Net+ [34], FISTA-Net [37], AMP-Net-K [39] and Nest-DGIL [41], as well as the plug-and-play image restoration method DPIR [45]. PD-Net stands out as a tried-and-true approach for dual-domain reconstruction. Likewise, other deep unfolding methods also perform corrections in both the image domain and the projection domain. The FBP with the ”Ram-Lak” filter, which was implemented with the Iradon transform [62], is utilized as the initial step for the proposed methods and other compared methods. The maximum number of iterations for FISTA-TV is set to 100, and the regularization parameter is tuned to its optimal value. As reported in [37], the number of stages for PD-Net, ISTA-Net, ISTA-Net+, FISTA-Net, AMP-Net-K, Nest-DGIL and our approach are all configured as $7$. A DRUNet model designed for sparse-view CT reconstruction is developed by changing the training data and replacing the noise level with the sparse-view down-sample scale. Meanwhile, the DPIR denoiser is re-trained using CT data. In DPIR, the data subproblem is tackled with the conjugate gradient (CG) method, which runs for five iterations. In addition, the outer iterations for DPIR are set to 40. Each of the compared methods is trained using an identical training dataset and employs the same techniques as described in their respective works.
@@ IV-C1 Quantitative evaluation
Table VI shows the average PSNR and SSIM values of the reconstruction results of the compared methods with different projection views of the downsampling. Clearly, the bottom row of Table VI shows the superior results. The DPIR denoiser trained on Gaussian noise cannot effectively remove the noise generated by sparse-view sampling. All of the DL methods and deep unfolding methods have somewhat competitive reconstruction results for fan-beam and parallel-beam projection data; however, our methods clearly outperform the other compared methods in all reconstruction tasks from different down-sampling projections. In particular for fan-beam projection, the proposed MVMS-RCN have a 4.49 dB improvement with 16 views, a 3.20 dB improvement with 32 views, and an average 1.96 dB improvement on five sparse view CT reconstruction tasks. Additionally, our methods have an average 0.39 dB improvement for parallel-beam projection.
In Fig. 2, we display the boxplots illustrating the Root Mean Square Error of Hounsfield Unit (RMSE (HU)) values for the fan-beam projection with 32 views and the parallel beam projection with 60 views to further highlight the superiority of our methods. Our methods are able to balance the trade-off between the use of prior knowledge from models and the powerful learning ability of DL, resulting in excellent generalizability. Furthermore, our methods outperform the compared methods in terms of reconstruction performance.
TABLE VII: Quantitative reconstruction comparisons (PSNR) of untrained sparse-view projections. The best and second best results are highlighted in Bold font and underlined ones, respectively.
@@
Figure 5: The convergence curve for the reconstruction of the perturbed projection data in a fan-beam projection with 32 views.
IV-C2 Qualitative evaluation
Fig. 3 shows the results of CT reconstruction using different methods in the fan-beam projection with 32 views. We can see that the FBP reconstruction has a lot of noise and artifacts. ISTA-Net, ISTA-Net+, FISTA-Net and Nest-DGIL can reduce some of the noise and streaking artifacts, but details are lost. RED-CNN, FBPConvNet and Uformer can remove streaking artifacts effectively, but some small structures may be smoothed out. PD-Net can remove streaking artifacts well and reconstruct high-quality CT images, but the details and texture information are not completely preserved. Although DPIR employs CG to maintain projection data consistency, it results in blurring of many fine details. Our results (l1)-(l4) have the best performance in eliminating noise artifacts while retaining intricate details (e.g., tiny blood vessels and bronchi).
@@ IV-D Blind Reconstruction of Untrained Sparse-view Projections
Table VII shows the quantitative comparison of blind reconstructions for untrained sparse-view projections. All compared methods, except for FBP, DRUNet, and DPIR, train separate models for each fan-beam projection task (with 16, 32, 64, 128, and 256 views) and parallel-beam projection tasks (having 30, 60, 90, 120, and 180 views). Conversely, the proposed framework utilizes a single model to handle five sparse-view CT reconstruction tasks. For methods that lack the capability to reconstruct CT images directly from untrained sparse-view projections, their PSNR results are evaluated using the two nearest trained sparse-view models.
Deep unfolding methods (PD-Net, ISTA-Net, ISTA-Net+ and Nest-DGIL) transform truncated unfolding optimization into an end-to-end trainable deep network. However, these methods typically require task-specific training, potentially reducing their effectiveness in practical applications. In contrast, our method removes the necessity for such specialized training, providing a more adaptable and easily implementable solution across various scenarios. This flexibility arises from our technique of refining projection errors from both projected and non-projected perspectives in both full-view and sparse-view environments, presenting a novel method to utilize available projection data. Furthermore, the CNN denoiser in DRUNet and DPIR lacks a comprehensive analysis and explanation. We convert the completion of details in the image domain into correcting errors of different scales related to the target image’s geometric errors across multi-scale and multi-resolution levels. By aligning the detailed mathematical framework with the network architecture, our approach ensures that all mathematical operations are systematically integrated into the network design. Consequently, the transition from mathematical theory to network design is more natural compared to existing deep unfolding methods.
Our framework outperforms the other methods for fan-beam projection, with an average improvement of 2.46 dB. Specifically, it improves at least 5.05 dB, 2.97 dB, 0.87 dB, and 0.97 dB for 24 views, 48 views, 96 views, and 192 views, respectively. For parallel-beam projection, our framework improves 1.28 dB, 0.21 dB, 0.16 dB, and 0.27 dB for 45 views, 75 views, 105 views, and 150 views, respectively, with an average improvement of 0.48 dB. It is noteworthy that our method can use the single multi-sparse-view model to produce more satisfactory results with any untrained sparse-view projections than the other compared methods.
@@ IV-E Plug-and-Play Sparse-view CT Reconstruction
Due to the shared multi-scale geometric correction module $\mathcal{D}$, the proposed MVMS-RCN can serve as a plug-and-play denoiser [45] for sparse-view CT reconstruction. To bridge the gap between simulated and real medical data, we introduced a perturbation of ±$\textpm 1\%$ to the distances between the rotation center and both the source and the detector, enhancing the realism of the medical CT reconstruction process. The plug-and-play sparse-view CT reconstruction results for the perturbed data, using fan-beam projection with 32 views, are shown in Fig. 5.
When encountering perturbed projection data, DPIR’s reconstruction performance unfortunately suffers from an undesirable decline due to over-fitting. However, MVMS-RCN demonstrates exceptional stability, with the reconstruction results continually improving with each iteration. Remarkably, MVMS-RCN achieves optimal performance after the 9th iteration, and the following iterations maintain consistent results without significant degradation, effectively minimizing the risk of over-fitting. This highlights the potential of MVMS-RCN as a flexible plug-and-play solution designed for sparse-view CT reconstruction. Unlike DPIR, which frequently demands manual parameter tuning, MVMS-RCN removes this requirement, significantly enhancing reproducibility and efficiency. The ability of MVMS-RCN to easily adapt and succeed in challenging scenarios involving perturbed data highlights its robustness and adaptability for practical applications, marking it as a promising advancement in the field of CT reconstruction.
Figure 6: The convergence curve for the reconstruction of the clinically rebinned projection data in a fan-beam projection with 36 views ($\times$32).
IV-F Unsupervised Fine-tuning for Clinical Reconstruction
In our clinical study concentrated on unsupervised fine-tuning, we utilize the same validation and test raw projection data from the ’Low Dose CT Image and Projection Data’ as detailed in Section IV-A. To improve computational efficiency, we adapt the different scanning configuration [66] for unsupervised training and evaluation, as detailed in Table 1 of [67], with a total of 1152 views and 736 detectors. The projection data for the validation and test patients are rebinned into 18, 36, 72, 144, and 288 views, with a voxel size of 0.7 in each slice. We select 362 slices from the validation patient and 385 slices from the test patient for use in unsupervised training and evaluation.
To adapt the pre-trained unified plug-and-play framework for a new scanning configuration, we minimize a loss function that operates without ground truth, defined by
$\begin{split}\mathcal{L}_{\text{total}}(\bm{\Theta})=&\frac{1}{d}\left\|\bm{P}_{s}^{T}\bm{P}_{s}\bm{x}_{n_{s}}-\bm{P}_{s}^{T}\bm{y}_{s}\right\|_{1}\\ &+\gamma\left(1-{\text{SSIM}}\left(\bm{P}_{s}^{T}\bm{P}_{s}\bm{x}_{n_{s}},\bm{P}_{s}^{T}\bm{y}_{s}\right)\right)\end{split}$
through unsupervised fine-tuning across 30 epochs using the validation dataset. We also use the no-reference image quality assessment metric known as the Natural Image Quality Evaluator (NIQE) [68] to assess plug-and-play clinically rebinned reconstruction, Fig.6 shows the convergence curve of the unsupervised sparse-view CT reconstruction derived from 36 views. The NIQE convergence trend closely aligns with that of PSNR, highlighting our method’s potential for unsupervised applications with clinical data.
@@ V Conclusion
In this study, we propose a novel interpretable dual-domain framework that can flexibly deal with multi-sparse-view CT imaging in a single model. The proposed framework integrates the theoretical advantages of model-based methods and the superior reconstruction performance of DL-based methods, providing the generalizability of DL. It gives us a new way to explore the potential projection information, offers a new perspective on designing an interpretable network, and serves as a flexible model for multi-sparse-view CT reconstruction. Experiments show that the proposed framework outperforms other state-of-the-art methods in terms of qualitative and quantitative evaluaions. In future work, our aim is to enhance the unified framework to simultaneously reconstruct fan-beam and parallel-beam projection data, and to extend its applicability to 3D reconstruction, limited-angle CT, and real medical CT image reconstruction problems. Our ultimate goal is to develop a more comprehensive and adaptable reconstruction framework.